OpenStudy (anonymous):

'Nother Q. Please Help. ;3

OpenStudy (anonymous):

@cvblondie712 @Ell @FlaCaLOL @Forever_Me @Savannah_Noelle @TSwizzle @ParthKohli @imbreanna

OpenStudy (ell):

Whats the question? (:

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

Oh yea...-.- wow sorry

OpenStudy (ell):

haha(: np

OpenStudy (ivettef365):

Just cross multiply to find value of x

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

5 7 ----=--- 2x-3 3x

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

(5)(3x) = 7(2x-3)

OpenStudy (ivettef365):

can you take it from there?

OpenStudy (anonymous):

Yes, probably.

OpenStudy (ivettef365):

What did you get?

OpenStudy (anonymous):

15x=14x-21

OpenStudy (ivettef365):

good and now just subtract 14x on each side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1x=-21

OpenStudy (anonymous):

x=-21

OpenStudy (anonymous):

ok now 27

OpenStudy (anonymous):

Suppose y varies inversly with x. write an equation for the inverse vaiation.

OpenStudy (anonymous):

y=7 when x=3

OpenStudy (anonymous):

@hartnn

OpenStudy (ivettef365):

Use formula y=kx

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that is..?

OpenStudy (anonymous):

hello?

OpenStudy (ivettef365):

Sorry that is the direct variation the inverse variation is y = k/x

OpenStudy (ivettef365):

which can also be read as yx= k

OpenStudy (ivettef365):

therefore you have (3)(7)=k

OpenStudy (ivettef365):

the inverse variation would be y = 21/k or you can write it as xy=21

OpenStudy (anonymous):

I know I figured it out same with 29